OpenStudy (anonymous):

Which inequality represents the solutions to a over five + 3 < 5? a < 10 a > 10 a < 40 a > 40

OpenStudy (whpalmer4):

\[\frac{a}{5} + 3 < 5\]You need to isolate \(a\), so subtract 3 from both sides, then multiply both sides by 5 to eliminate the fraction and leave you with \(a\) on the left. What do you get?

OpenStudy (anonymous):

a<10 a>10 a<40 a>40

OpenStudy (anonymous):

one of these is the answer

OpenStudy (whpalmer4):

The one that you get if you do the steps I gave you.

jimthompson5910 (jim_thompson5910):

The answer will depend on if the inequality is \[\Large \frac{a}{5} + 3 < 5\] OR if it's \[\Large \frac{a}{5+3} < 5\] Most likely it is probably the first one (since the 5+3 can easily be replaced with 8), but it's possible that they actually wrote "5+3" instead of "8"

OpenStudy (whpalmer4):

I took it as the first one because she wrote "a over five" as words, then "+3"

OpenStudy (anonymous):

is someone gives me the answer then i will give them a medal

OpenStudy (whpalmer4):

we've got thousands of medals. How about you work out the answer from what we told you?

OpenStudy (whpalmer4):

I'll give you a medal if you do.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

a<10

OpenStudy (anonymous):

i dont know

jimthompson5910 (jim_thompson5910):

Assuming the inequality is \[\Large \frac{a}{5} + 3 < 5\] we must subtract 3 from both sides. Afterward, we multiply both sides by 5 to completely isolate 'a' \[\Large \frac{a}{5} + 3 < 5\] \[\Large \frac{a}{5} < 5-3\] \[\Large \frac{a}{5} < 2\] \[\Large a < 2*5\] \[\Large a < 10\] So you are correct.

jimthompson5910 (jim_thompson5910):

Solving inequalities is just like solving equations

OpenStudy (anonymous):

yes you get a medal

jimthompson5910 (jim_thompson5910):

The only difference is that the inequality sign will flip when multiplying/dividing both sides by a negative number

OpenStudy (anonymous):

thank you so much *hug*

jimthompson5910 (jim_thompson5910):

yw